OpenStudy (anonymous):

Which is a prime number? A. 21 B. 31 C. 57 D. 91

OpenStudy (anonymous):

which number can only be divided by 1 and itself?

OpenStudy (anonymous):

D,B,C

OpenStudy (eyust707):

7 * 3 = 21... not prime 19 * 3 = 57.. not prime 9 * 9 = 81.. not prime